# Ferraz, Cappart, and Vidal

### v1

@@ 4 Graph Neural Network for District-Cost Estimations
We focus on the estimation of delivery costs for the districts. A direct calculation of Equation (2) by solving TSPs on different demand scenarios is fast enough to evaluate the district costs of any given solution of the districting-and-routing problem. However, it is impractically slow within a search method for the districting problem (e.g., a local search) due to the large number of candidate districts considered through the search for move evaluations. To provide fast and accurate estimations for such applications, we explore the option to learn an approximation of the delivery costs by supervised learning, more specifically by relying on the considerable methodological progress recently made on deep learning and graph neural networks (GNNs). This approach is described in the remainder of the section, starting from the feature information used for training, the architecture of the network, and the training process.
@@ 4.1 Features
Our GNN leverages the same undirected graph $G(V,E)$ as the one depicted in Section 3, with the vertices $V$ corresponding to BUs and edges in $E$ correspond to adjacency relations. For learning the GNN weights, we rely on a training set $\mathcal{D}^{\textsc{Train}}$ of districts. Each district $d\in\mathcal{D}^{\textsc{Train}}$ is characterized by a vector of features on each node of the graph and an estimated cost value using SAA as in Equation (2) with $500$ previous demand realization scenarios for each BU. This estimated cost, also referred to as the label, is the value we want to predict with the GNN. We use eight features $\bm{f}_{vd}=(\xi_{v},\sqrt{\xi_{v}},q_{v},a_{v},\sqrt{a_{v}},\rho_{v},\delta_{v},e_{vd})^{\top}\in\mathbb{R}^{8}$ for each vertex $v$ (associated to a BU) and each district $d$ in the training set:
@@
    an inclusion variable $e_{vd}$, taking value $1$ if BU $v$ belongs to $d$ and 00 otherwise.
It is noteworthy that the first seven features remain fixed when considering examples generated in the same metropolitan area. Only the last feature changes, according to the current district considered in the sample.
@@ 4.2 Architecture of the graph neural network
    a node embedding layer, taking as input a district sample in the format described previously and whose final output is a latent vector of features for each node of the graph. These outputs are also referred to as node embedding. Intuitively, it is a function that aggregates feature information defined over a graph into a vectorial representation of the features by aggregating information from neighboring nodes. This aggregation is done several times and corresponds to a layer of the GNN.
@@
Node Embedding Layers. Let $G(V,E)$ be the graph representation of a metropolitan area, and $\bm{f}_{vd}$ be the features of each node $v\in V$ for a district $d$. Formally, a GNN computes a $m$-dimensional features embedding $\bm{\mu}_{v}\in\mathbb{R}^{m}$ for each node $v\in V$ in $G$ (i.e., the node embedding). The node features $\bm{f}_{vd}$ are aggregated iteratively with the neighboring nodes in the graph. After a predefined number of aggregation steps, the embedding of each node is produced and encompasses both local and global characteristics of the graph. These operations can be carried out in different ways (Dai, Dai, and Song 2016, Veličković et al. 2018, Hamilton, Ying, and Leskovec 2017), and public implementations are available for many of these architectures (Wang 2019, Fey and Lenssen 2019). Following Khalil et al. (2017) who considered graph neural networks solving combinatorial problems over graphs, this paper is based on their implementation referred to as Structure2Vec (Dai, Dai, and Song 2016). We note that other architectures based on GNNs could be used as well.
@@
The idea is to compute the new embedding $\bm{\mu}_{v}^{t+1}$ as a parametrized sum of the previous embedding ($\bm{\mu}_{v}^{t+1}$) and the node features ($\bm{f}_{vd}$). Following the recommendations of the initial implementation, four aggregation steps are done ($T=4$), each hidden embedding is a vector of $64$ values ($\bm{\mu}_{v}^{2},\bm{\mu}_{v}^{3},\bm{\mu}_{v}^{4}\in\mathbb{R}^{64}$). Then, a last aggregation step is performed on the last hidden embedding. This is presented in Equation (8), where $\bm{\theta}_{3}\in\mathbb{R}^{k\times p}$ is another weight tensor. In our case, $k$ is set to 1024, which yields a $1024$-dimensional vector as the output embedding ($\bm{\mu}_{v}^{\texttt{out}}\in\mathbb{R}^{1024}$):
@@
Graph Embedding layer. Once a vectorized representation $\bm{\mu}_{v}^{\texttt{out}}$ has been computed for each vertex $v$, this information is used to compute $\bm{z}$, a vectorized representation of the entire graph. This is done by summing together the embedding of each node and applying a non-linear transformation (e.g., ReLU) to the result. This is illustrated in Equation (9) where $\bm{\theta}_{4}\in\mathbb{R}^{k\times k}$ and $\bm{\theta}_{5}\in\mathbb{R}^{k\times k}$ are two other tensors of weights that must be learned ($k=1024$). This is also referred to as a pooling operation:
$\bm{z}=\bm{\theta}_{4}\texttt{ReLU}\Big{(}\bm{\theta}_{5}\sum_{v\in V}\bm{\mu}_{v}^{\texttt{out}}\Big{)}.$ (9)
@@
Fully-Connected Neural Network. Finally, the embedding $\bm{z}$ goes through a standard fully connected neural network of $2$ layers, with $100$ neurons for the hidden layer and a single neuron for the output layer. This neural network can be represented as a function $\texttt{FCNN}:\mathbb{R}^{1024}\to\mathbb{R}$, that computes the expected delivery cost $\hat{\Phi}(d)$ inside a district $d$ thanks to the pre-computed graph embedding and two additional tensor of weights: $\bm{\theta}_{6}\in\mathbb{R}^{100\times 1024}$ and $\bm{\theta}_{7}\in\mathbb{R}^{1\times 100}$. Assembling all the pieces together, the delivery cost of a district $d$, represented as a graph $G$, and the features $\bm{f}_{vd}$ for each $v\in V$ are computed as follows:
$\hat{\Phi}(d)=\texttt{FCNN}\Bigg{(}\texttt{GNN}\Big{(}G,\big{[}\bm{f}_{vd}~{}\big{|}~{}v\in V\big{]}~{};~{}\bm{\theta}_{1},\dots,\bm{\theta}_{5}\Big{)}~{};~{}\bm{\theta}_{6},\bm{\theta}_{7}\Bigg{)}.$ (10)
@@ 4.3 Training
The network is trained using back-propagation using the mean absolute error as loss function (L1 loss) to find good values of the tensors $\bm{\theta}_{1}$, $\bm{\theta}_{2}$, $\bm{\theta}_{3}$ $\bm{\theta}_{4}$, $\bm{\theta}_{5}$, $\bm{\theta}_{6}$, and $\bm{\theta}_{7}$. The training is carried out for a maximum of 24 hours or $10,000$ epochs (with a batch size of 64) using Adam optimizer (Kingma and Ba 2014). As an additional early stopping criterion, the training is aborted when no reduction in the loss is observed after $1,000$ consecutive epochs. We used the default values for the optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$, and no weight decay), except for the learning rate, which was set to $10^{-4}$ instead of $10^{-3}$ to better stabilize the training.
@@ 5 Districting Solution Strategy
The trained GNN provides us with a fast estimation oracle for district costs. With this, we can now focus our attention on the solution to the districting-and-routing problem. As seen in Section 3, solving the problem through its graph-partitioning mathematical formulation is impractical for most medium- and large-scale instances, due to the difficulty of modeling the objective and the large number of possible districts. To permit experimentation on larger cases, we rely on a fairly standard local search-based approach, which only requires access to the district-cost evaluation oracle. As seen in our numerical experiments and in Appendix 8, this method closely matches the results of the exact set-partitioning-based algorithm on small cases solvable to optimality and provides a faster and more scalable solution-approach substitute otherwise. Our approach is built upon the Iterated Local Search (ILS) principle (Lourenço, Martin, and Stützle 2019), a well-known strategy to guide local search heuristics beyond local optima. As summarized in Algorithm 1, it simply consists of iteratively applying a local search heuristic from an initial solution to attain a local minimum and then perturbing this solution to generate a new starting point for the local search. This process is iterated until a stopping criterion. The section briefly describes the different components of the ILS: the construction of an initial solution, the local search moves, and the perturbation scheme.
@@
2       $s\leftarrow\text{Pertubation}(s)$ $s\leftarrow\text{Local Search}(s)$ if $\text{\emph{cost}}(s)<\text{\emph{cost}}(s_{\textsc{Best}})$ then $s_{\textsc{Best}}\leftarrow s$
return $s_{\textsc{Best}}$
@@ 5.1 Initial Solution
Due to the cardinality constraints on the number of BUs per district (between $n_{\textsc{l}}$ and $n_{\textsc{u}}$) and the requirement that the districts must remain connected, even locating an initial feasible solution of the districting-and-routing problem is NP-complete by reduction from the balanced connected $k$-partition problem (BCP-$k$Dyer and Frieze 1985). Consequently, we rely on the solution of a mixed integer linear program to construct an initial feasible solution, using a compact network-flow formulation of the BCP-$k$ from Miyazawa et al. (2020).
Given the initial graph $G(V,E)$ representing the adjacency between the different BUs, the network flow formulation uses an auxiliary directed graph $H(V\cup S,A)$ in which $S$ is a set of $k$ source vertices, one for each district. The set of arcs $A$ is obtained by including two opposite arcs $(u,v)$ and $(v,u)$ for each edge $\{u,v\}\in E$, as well as an arc between each $u\in S$ and $v\in V$. Binary design variables $y_{ij}$ take value $1$ when $(i,j)$ carries some flow and $f_{ij}$ is the amount of flow. With this, the feasibility model seeks a solution that satisfies:
$\sum_{u\in V\cup S}f_{uv}-\sum_{u\in V\cup S}f_{vu}=1$ $\forall v\in V$ (11)
@@
This model is constructed in a way that sources $s\in S$ representing districts send flow towards vertices $v\in V$ representing the BUs. Constraints (11) ensure flow conservation and guarantee that each BU vertex $v\in V$ receives one unit of flow. Constraint (12) sets a limit on the maximum flow of an arc and consequently enforces the upper limit $n_{\textsc{u}}$ on the number of BUs per district. Constraint (13) ensures that each source vertex feeds flow to at least $n_{\textsc{l}}$ BUs. Constraint (14) orders the sources by flow amount to eliminate symmetry. Finally, Constraints (1516) uniquely match each source vertex of $S$ with a single receiving BU vertex of $V$ (which will then forward flow to other connected BU vertices – ensuring the connectivity constraints) and ensure that each BU is connected to a single BU or source. Any feasible solution of this model can be converted into an initial solution of the districting-and-routing problem, by assigning each BU vertex $v\in V$ to the (unique) district vertex $s\in S$ from which its flow originated.
@@ 5.2 Local Search
After the initial solution construction and after each perturbation, our solution approach applies a local search (LS) procedure. As summarized in Algorithm 2, this LS follows a first improvement policy that consists of exploring a neighborhood in a random order and directly applying any improving move until a local minimum is attained. For any solution $s$, the neighborhood $\mathcal{N}(s)$ is defined as the set of solutions that can be attained by applying any of these two moves: (1) Relocate$(u,d)$, which reassigns a BU $u$ to a new district $d$, and (2) Swap$(u,v)$, which exchanges the districts of two BUs $u$ and $v$. Therefore, the neighborhood size is such that $|\mathcal{N}(s)|=\mathcal{O}(kn+n^{2})=\mathcal{O}(n^{2})$, but moves leading to infeasible solutions w.r.t. district-cardinality and connectivity constraints can be directly disregarded. This permits several speedup strategies. Firstly, moves can be limited to pairs of districts $d_{i}$ and $d_{j}$ that are connected, such that there exists two neighbors BUs $(u,v)\in E$ with $u\in d_{i}$ and $v\in d_{j}$. Besides this, we can define border $\mathcal{B}_{ij}$ as the set of BUs of a district $d_{i}$ that have at least one neighbor BU belonging to district $d_{j}$. We can then further limit the set of moves between districts $d_{i}$ and $d_{j}$ to BUs $u\in\mathcal{B}_{ij}$ and $v\in\mathcal{B}_{ji}$.
@@
3             Find and apply, if improving, the best move on $s$ among:     $\bullet$ all feasible relocate moves for $\{(u,d_{j})\,|\,u\in\mathcal{B}_{ij}\}\cup\{(v,d_{i})\,|\,v\in\mathcal{B}_{ji}\}$;     $\bullet$ all feasible swap moves for $\{(u,v)\,|\,u\in\mathcal{B}_{ij},\,v\in\mathcal{B}_{ji}\}$;
@@ 5.3 Perturbation
The perturbation operator is an important building block of any ILS. It must permit escaping from local optima without losing too much structural information on good solutions. Our perturbation procedure simply uses the same set of moves as the local search but applies any feasible move with a probability $p_{\textsc{rm}}$ regardless of their cost. Parameter $p_{\textsc{rm}}$ drives the amplitude of the perturbation. Its calibration is discussed in Appendix 8.
6 Numerical Experiments
For our experimental analyses, we focus on districting-and-routing problem instances that occur in five metropolitan areas in the UK (London, Bristol, Manchester, Leeds, and West-Midlands) with very diverse geographical characteristics. The goal of our experiments is twofold. First, we aim to evaluate how accurate the proposed GNN is in terms of routing-cost estimation accuracy. Next, we measure to which extent cost-estimation accuracy impacts the ability to make good strategic decisions, i.e., to what extent solutions of districting problems using the GNN or other cost predictors differ in terms of their partitions and operational efficiency.
In the rest of this section, we discuss the data collection and generation of test instances (Section 6.1), as well as the baseline methods considered for routing-cost estimation (Section 6.2). Then, we analyze the accuracy of the different prediction cost-estimation methods (Section 6.3), and we measure the impact of different estimation algorithms when optimizing districts as well as the characteristics (e.g., compactness) of the partitions thereby generated (Section 6.4).

### v2

@@ 4 Graph Neural Network for District-Cost Estimations
We focus on the estimation of delivery costs for the districts. A direct calculation of Equation (2) by solving TSPs on different demand scenarios is fast enough to evaluate the district costs of any given solution of the districting-and-routing problem. However, it is impractically slow within a search method for the districting problem (e.g., a local search) due to the large number of candidate districts considered through the search for move evaluations. To provide fast and accurate estimates for such applications, we explore the option to learn an approximation of the delivery costs by supervised learning, more specifically by relying on the considerable methodological progress recently made on deep learning and graph neural networks (GNNs). This approach is described in the remainder of the section, starting from the feature information used for training, the architecture of the network, and the training process.
@@ 4.1 Features
Our GNN leverages the same undirected graph $G(V,E)$ as the one depicted in Section 3, with the vertices $V$ corresponding to BUs and edges in $E$ corresponding to adjacency relations. For learning the GNN weights, we rely on a training set $\mathcal{D}^{\textsc{Train}}$ of districts. Each district $d\in\mathcal{D}^{\textsc{Train}}$ is characterized by a vector of features on each node of the graph and an estimated cost value using SAA as in Equation (2) with $500$ previous demand realization scenarios for each BU. This estimated cost, also referred to as the label, is the value we want the GNN to predict. We use eight features $\bm{f}_{vd}=(\xi_{v},\sqrt{\xi_{v}},q_{v},a_{v},\sqrt{a_{v}},\rho_{v},\delta_{v},e_{vd})^{\top}\in\mathbb{R}^{8}$ for each vertex $v$ (associated to a BU) and each district $d$ in the training set. The features used are:
@@
    a membership variable $e_{vd}$, taking value $1$ if BU $v$ belongs to $d$ and 0 otherwise.
The first seven (continuous) features were normalized to $[0,1]$ by dividing by their greatest value. It is noteworthy that they remain fixed when considering examples generated in the same metropolitan area. Only the last feature changes, according to the current district considered in the sample. The choice of these features follows well-established CA formulas, notably BHHD, FIG, and SNN (see Section 2). Population and its square root, area and its square root, and the average distance to the depot are standard descriptors in these CA models, to which we additionally include the BU perimeter. Together, these variables summarize the main topological and geographical characteristics of the BUs.
@@ 4.2 Architecture of the graph neural network
    a node embedding layer, taking as input a district sample in the format described previously and whose final output is a latent vector of features for each node of the graph. These outputs are also referred to as node embedding. Intuitively, it is a function that converts feature information defined over a graph into a vector representation of the features by aggregating information from neighboring nodes. This aggregation is done several times and corresponds to a layer of the GNN.
@@
Node Embedding Layers. Let $G(V,E)$ be the graph representation of the metropolitan area, and $\bm{f}_{vd}$ be the features of each node $v\in V$ for a district $d$. Formally, a GNN computes a $m$-dimensional features embedding $\bm{\mu}_{v}\in\mathbb{R}^{m}$ for each node $v\in V$ in $G$ (i.e., the node embedding). The node features $\bm{f}_{vd}$ are aggregated iteratively with the neighboring nodes in the graph. After a predefined number of aggregation steps, the embedding of each node is produced and captures both local and global characteristics of the graph. These operations can be carried out in different ways (Dai, Dai, and Song 2016, Veličković et al. 2018, Hamilton, Ying, and Leskovec 2017), and public implementations are available for many of these architectures (Wang 2019, Fey and Lenssen 2019). Following Dai et al. (2017), who considered graph neural networks solving combinatorial problems over graphs, this paper is based on their implementation referred to as Structure2Vec (Dai, Dai, and Song 2016). We note that other architectures based on GNNs could be used as well.
@@
The idea is to compute the new embedding $\bm{\mu}_{v}^{t+1}$ as a parametrized sum of the previous embedding ($\bm{\mu}_{v}^{t}$) and the node features ($\bm{f}_{vd}$). Following the recommendations of the initial implementation, four aggregation steps are done ($T=4$), each hidden embedding is a vector of $64$ values ($\bm{\mu}_{v}^{2},\bm{\mu}_{v}^{3},\bm{\mu}_{v}^{4}\in\mathbb{R}^{64}$). Then, a last aggregation step is performed on the last hidden embedding. This is presented in Equation (8), where $\bm{\theta}_{3}\in\mathbb{R}^{k\times p}$ is another weight tensor. In our case, $k$ is set to 1024, which yields a $1024$-dimensional vector as the output embedding ($\bm{\mu}_{v}^{\texttt{out}}\in\mathbb{R}^{1024}$):
@@
Graph Embedding layer. Once a vector representation $\bm{\mu}_{v}^{\texttt{out}}$ has been computed for each vertex $v$, this information is used to compute $\bm{z}$, a vector representation of the entire graph. This is done by summing together the embedding of each node and applying a non-linear transformation (e.g., ReLU) to the result. This is also referred to as a pooling operation:
$\bm{z}=\texttt{ReLU}\Big(\sum_{v\in V}\bm{\mu}_{v}^{\texttt{out}}\Big).$ (9)
@@
Fully-Connected Neural Network. Finally, the embedding $\bm{z}$ goes through a standard fully connected neural network of two layers, with $100$ neurons for the hidden layer and a single neuron for the output layer. This neural network can be represented as a function $\texttt{FCNN}:\mathbb{R}^{1024}\to\mathbb{R}$, that computes the expected delivery cost $\hat{\Phi}(d)$ inside a district $d$ thanks to the pre-computed graph embedding and two additional tensor of weights: $\bm{\theta}_{4}\in\mathbb{R}^{100\times 1024}$ and $\bm{\theta}_{5}\in\mathbb{R}^{1\times 100}$. Assembling all the pieces together, the delivery cost of a district $d$, represented as a graph $G$, and the features $\bm{f}_{vd}$ for each $v\in V$ are computed as follows:
$\hat{\Phi}(d)=\texttt{FCNN}\Bigg(\texttt{GNN}\Big(G,\big[\bm{f}_{vd}~\big|~v\in V\big]~;~\bm{\theta}_{1},\dots,\bm{\theta}_{3}\Big)~;~\bm{\theta}_{4},\bm{\theta}_{5}\Bigg).$ (10)
@@ 4.3 Training
The network, parameterized by all $\theta$ tensors (171,465 free parameters), is trained via back-propagation using the mean absolute error as loss function (L1 loss). The training is carried out for a maximum of 24 hours or $10,000$ epochs (with a batch size of 64) using Adam optimizer (Kingma and Ba 2014). As an additional early stopping criterion, the training is aborted when no reduction in the loss is observed after $1,000$ consecutive epochs. We used the default values for the optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$, and no weight decay), except for the learning rate, which was set to $10^{-4}$ instead of $10^{-3}$ to better stabilize the training.
@@ 5 Districting Solution Strategy
The trained GNN provides us with a fast estimation oracle for district costs. With this, we can now focus our attention on the solution to the districting-and-routing problem. As seen in Section 3, solving the problem through its graph-partitioning mathematical formulation is impractical for most medium- and large-scale instances, due to the difficulty of modeling the objective and the large number of possible districts. To permit experimentation on larger cases, we rely on a fairly standard local search-based approach, which only requires access to the district-cost evaluation oracle. As seen in our numerical experiments and in the Appendix, this method closely matches the results of the exact set-partitioning-based algorithm on small cases solvable to optimality and provides a faster and more scalable solution approach otherwise. Our approach is built upon the Iterated Local Search (ILS) principle (Lourenço, Martin, and Stützle 2019), a well-known strategy for guiding local search heuristics beyond local optima. As summarized in Algorithm 1, it simply consists of iteratively applying a local search heuristic from an initial solution to attain a local minimum and then perturbing this solution to generate a new starting point for the local search. This process is iterated until a stopping criterion is met. The section briefly describes the different components of the ILS: the construction of an initial solution, the local search moves, and the perturbation scheme.
@@
2 $s\leftarrow\text{Local Search}(s)$
3 $s_{\textsc{Best}}\leftarrow s$
4 while termination criteria not attained
5 $s\leftarrow\text{Pertubation}(s)$
6 $s\leftarrow\text{Local Search}(s)$
7
8 if $\text{\emph{cost}}(s)<\text{\emph{cost}}(s_{\textsc{Best}})$ then $s_{\textsc{Best}}\leftarrow s$
9
10return $s_{\textsc{Best}}$
@@ 5.1 Initial Solution
As discussed in Section 3, finding an initial solution of the districting-and-routing problem is NP-hard due to the connectivity and district balance constraints. Consequently, we rely on the solution of an MILP to construct an initial feasible solution, using a compact network-flow formulation of the balanced connected $k$-partition problem from Miyazawa et al. (2020).
Given the initial graph $G(V,E)$ representing the adjacency between the different BUs, the network flow formulation uses an auxiliary directed graph $H(V\cup S,A)$ in which $S$ is a set of $k$ source vertices, one for each district. The set of arcs $A$ is obtained by including two opposite arcs $(u,v)$ and $(v,u)$ for each edge $\{u,v\}\in E$, as well as an arc between each $u\in S$ and $v\in V$. Binary design variables $y_{ij}$ take the value $1$ when $(i,j)$ carries some flow, and $f_{ij}$ is the amount of flow. With this, the feasibility model seeks a solution that satisfies:
$\hskip-5.69046pt\sum_{u\in V\cup S}f_{uv}-\sum_{u\in V\cup S}f_{vu}=1$ $\forall v\in V$ (11)
@@
This model is constructed so that sources $s\in S$ representing districts send flow towards vertices $v\in V$ representing the BUs. Constraints (11) ensure flow conservation and guarantee that each BU vertex $v\in V$ receives one unit of flow. Constraint (12) sets a limit on the maximum flow of an arc and consequently enforces the upper limit $n_{\textsc{u}}$ on the number of BUs per district. Constraint (13) ensures that each source vertex feeds flow to at least $n_{\textsc{l}}$ BUs. Constraint (14) orders the sources by flow amount to eliminate symmetry. Finally, Constraints (1516) uniquely match each source vertex of $S$ with a single receiving BU vertex of $V$ (which will then forward flow to other connected BU vertices – ensuring the connectivity constraints) and ensure that each BU is connected to a single BU or source. Solving this MILP on graphs with a few hundred BUs (as in our cases) takes seconds using modern solvers (Gurobi v12.0.0). Then, any feasible solution of this model can be converted into an initial solution of the districting-and-routing problem by assigning each BU vertex $v\in V$ to the (unique) district vertex $s\in S$ from which its flow originated.
@@ 5.2 Local Search
After the initial solution construction and after each perturbation, our solution approach applies a local search (LS) procedure. As summarized in Algorithm 2, this LS follows a first-improvement policy that consists of exploring a neighborhood in a random order and directly applying any improving move until a local minimum is attained. For any solution $s$, the neighborhood $\mathcal{N}(s)$ is defined as the set of solutions that can be attained by applying any of these two moves: (1) Relocate(u,d)(u,d), which reassigns a BU $u$ to a new district $d$, and (2) Swap(u,v)(u,v), which exchanges the districts of two BUs $u$ and $v$. Therefore, the neighborhood size is such that $|\mathcal{N}(s)|=\mathcal{O}(kn+n^{2})=\mathcal{O}(n^{2})$, but moves leading to infeasible solutions w.r.t. district-cardinality and connectivity constraints can be directly disregarded. This permits several speedup strategies. Firstly, moves can be limited to pairs of districts $d_{i}$ and $d_{j}$ that are connected, such that there exist two neighbors BUs $(u,v)\in E$ with $u\in d_{i}$ and $v\in d_{j}$. Besides this, we can define border $\mathcal{B}_{ij}$ as the set of BUs of a district $d_{i}$ that have at least one neighbor BU belonging to district $d_{j}$. We can then further limit the set of moves between districts $d_{i}$ and $d_{j}$ to BUs $u\in\mathcal{B}_{ij}$ and $v\in\mathcal{B}_{ji}$.
@@
4      Find and apply, if improving, the best move on $s$ among:
5       $\bullet$ all feasible relocate moves for $\{(u,d_{j})\,|\,u\in\mathcal{B}_{ij}\}\cup\{(v,d_{i})\,|\,v\in\mathcal{B}_{ji}\}$;
6       $\bullet$ all feasible swap moves for $\{(u,v)\,|\,u\in\mathcal{B}_{ij},\,v\in\mathcal{B}_{ji}\}$;
@@ 5.3 Perturbation
The perturbation operator is an important building block of any ILS. It must permit escaping from local optima without losing too much structural information on good solutions. Our perturbation procedure simply uses the same set of moves as the local search but applies any feasible move with a probability $p_{\textsc{rm}}$ regardless of its cost. Parameter $p_{\textsc{rm}}$ drives the amplitude of the perturbation. Its calibration is discussed in the Appendix.
6 Experimental Analyses
For our experimental analyses, we focus on districting-and-routing problem instances that occur in five metropolitan areas in the UK (London, Bristol, Manchester, Leeds, and the West Midlands) with very diverse geographical characteristics. The goal of our experiments is twofold. First, we aim to evaluate the proposed GNN’s routing-cost estimation accuracy. Next, we measure the extent to which cost-estimation accuracy impacts the ability to make good strategic decisions, i.e., to what extent solutions of districting problems using the GNN or other cost predictors differ in terms of their partitions and operational efficiency.
In the rest of this section, we discuss the data collection and generation of test instances (Section 6.1), as well as the baseline methods considered for routing-cost estimation (Section 6.2). Then, we analyze the accuracy of the different prediction cost-estimation methods (Section 6.3), and measure the impact of different estimation algorithms when optimizing districts as well as the characteristics (e.g., compactness) of the partitions thereby generated (Section 6.4).